OpenStudy (anonymous):

maybe someone can help me better then my professor can.: why is this function continuous at every number in its domain. F(x) = 2x^2 − x − 2/ x^2 + 4 State the domain.

OpenStudy (anonymous):

when i found the domain i ended up with x-1/x+2 why is this wrong

OpenStudy (anonymous):

Because the numerator is continuous as it is a polynomial and the denominator is also a polynomial that has no zeroes, then the domain is:\[]-\infty,+\infty[\]

OpenStudy (anonymous):

okay that makes sense, so what did i do lol

OpenStudy (anonymous):

tangle with it?

OpenStudy (anonymous):

probably so, Thank You much

OpenStudy (anonymous):

when you have a fraction, always look at the denominator first. If there are zeroes, then it is not continuous

OpenStudy (anonymous):

so maybe you can help me with this problem also: For what value of the constant c is the function f continuous on (−infinity, infinity)? f(x) = cx2 + 5x if x < 6 x3 − cx if x ≥ 6 I dont understand the steps and finding this answer . the whole infinity thing throws me off

OpenStudy (anonymous):

Please, close the previous post and create a new one for your new question. Otherwise you cannot award medals again